Falconmaster:

A planet has an average distance to the sun of 4.54 AU. In two or more complete sentences, explain how to calculate the orbital period of the planet and calculate it.

Falconmaster:

@Shadow

Falconmaster:

@Vocaloid

Falconmaster:

help?

563blackghost:

I do believe you would use Kepler's Law. \(\large \bf{T_{p}=T_{e}\times (\frac{R_{p}}{R_{e}})^{\frac{3}{2}}}\) \(\bf{T_{e}=~orbital~period}\) Which is... \(\large\bf{T_{p}=1 \times (\frac{4.54}{1}})^{\frac{2}{3}} \approx 9.67~ years\)